# Wavespeed selection of travelling wave solutions of a two-component reaction-diffusion model of cell invasion

### v1

@@ 3 Calculation of travelling wave solutions
In this section we will compute the travelling wave solutions to (5). Let $z=x-ct$ be the travelling wave coordinate, and let $u(x,t)=U(z)$, $v(x,t)=V(z)$, and $u_{x}(x,t)=U^{\prime}(z)=W(z)$. The system of PDEs then becomes the system of three autonomous differential equations:
$U^{\prime}$ $=W$ (9a)
$V^{\prime}$ $=\frac{\gamma}{c}UV$ (9b)
$W^{\prime}$ $=\frac{1}{1-V}\left(\frac{\gamma}{c}UVW-cW-U(1-U-V)\right).$ (9c)
@@
The eigenvalues and eigenvectors of $(1,0,0)$ are
@@
$\lambda_{1}^{\prime}=0,\quad\lambda_{2}^{\prime}=-\mathcal{C}+\sqrt{\mathcal{C}^{2}-1},\qquad\lambda_{3}^{\prime}=-\mathcal{C}-\sqrt{\mathcal{C}^{2}-1},\quad\mathcal{C}=\frac{c}{2(1-V_{\infty})},$ (11)
@@
The zero eigenvalue $\lambda_{1}^{\prime}$ (with eigenvector pointing along the $V$-axis) is due to the non-isolated nature of each fixed point. The other two eigenvalues are negative, indicating that trajectories are attracted to the $V$-axis in its neighbourhood. As there is one degree of freedom in trajectories leaving $(1,0,0)$, there is a one-parameter family of trajectories that connect $(1,0,0)$ to the $V$-axis, each ending at a different value of $V_{\infty}$ between 00 and $1$.
In Colson et al. [7] trajectories of their similar system were found via a shooting algorithm. Here we compute solutions of (9) numerically by solving it as a boundary value problem on a large domain $z\in(0,z_{\infty})$, using the numerical continuation package Auto-07p [9]. This will allow us to compute solutions for more extreme values of parameters (in particular $\gamma$). Appropriate boundary conditions are
$U(0)=1+\varepsilon,\quad W(0)=-\epsilon\lambda_{2},\quad V(z_{\infty})=V_{\infty}$
where $\varepsilon$ is a small parameter, and $z_{\infty}\gg 1$. These boundary conditions enforce that near $(1,0,0)$, trajectories are on the eigenvector with the smaller positive eigenvalue. For a given $\gamma$ and $c$ we are able to construct travelling wave solutions for any value of $V_{\infty}$ between 0 and 1, by starting at $V_{\infty}=0$ (at which point $V$ vanishes identically, and $U,W$ are solutions to the Fisher–KPP system), and then increasing $V_{\infty}$ by numerical continuation.
We plot examples of such trajectories, projected onto the $(U,V)$ plane, in Fig. 4. In this figure we choose two sets of parameters $(\gamma,c)=(1,1)$ and $(10,1.24)$ in order to compare to the travelling wave solutions observed in the PDE simulations shown in Fig. 1.
@@
For given values of $\gamma$ and $c$, there is a one-parameter family of travelling wave solutions, that we may consider as characterised by the value $V_{\infty}$. Since the wave speed $c$ is a parameter of the system, we thus expect that for a given $V_{\infty}$ there is a one-parameter family of travelling wave solutions, parametrised by $c$, and it is not yet apparent which travelling wave solution will correspond to the long-time behaviour of a solution to the PDE system (5) for a given initial condition. In order to determine this wavespeed selection, we need to consider the behaviour of trajectories near the front of the travelling wave, that is, near the $V$-axis. Firstly, we note that for a given wavespeed $c$, there is a critical value $V=V_{c}$, at which the eigenvalues $\lambda^{\prime}_{2,3}$ switch from complex to real:
@@
This value is independent of the decay rate $\gamma$. Trajectories for which $V_{\infty}<V_{c}$ spiral into the relevant point on the $V$-axis, and since this corresponds to negative values of $U$, such trajectories cannot correspond to solutions to (5) that have evolved from a non-negative initial condition. This property is determined locally, and is equivalent to the condition used to determine the minimum wave speed in the Fisher–KPP equation.
However, it is not always the case that $U$ is positive on trajectories for which $V_{\infty}>V_{c}$. This is true in particular when $\gamma$ is sufficiently large. In Fig. 4(b) (in particular the inset of that figure) in which $\gamma=10$, we observe that trajectories for which $V_{\infty}$ is slightly larger than $V_{c}$ still approach the $V$-axis from the negative $(U<0)$ side, despite the fact that eigenvalues are real at these points. In this parameter regime there is a point $V_{s}$ (indicated as a red dot in Fig. 4(b)) that separates trajectories that have $U$ become negative, from those that remain positive. The existence of such a threshold value was also noted in the similar model of Colson et al. [7]. Unlike the point $V_{c}$, the point $V_{s}$ is not determined as a local property of the phase space (it is a property of the manner in which the unstable manifold of $(1,0,0)$ intersects the $V$-axis), so is unlikely to have a closed-form expression.
@@
Trajectories that end at $V_{\infty}>\max\{V_{c},V_{s}\}$ are also valid travelling wave solutions that exhibit a specific exponential decay. Indeed, when we plot the curves in the $(U,V)$ plane that correspond to the late-time PDE simulations (see Fig. 4), we see that these exactly correspond to trajectories on the phase plane, terminating either at $V_{c}$ or $V_{s}$ (for the compactly supported initial conditions), or at a higher value (for the slowly exponentially decaying initial condition). Identifying the eigenvalue $\lambda_{2}^{\prime}$ from (11) with $-a$, a travelling wave solution will have decay rate $a$ given that the wave speed (8) holds (which is how that formula is derived). However, as observed in the PDE simulations depicted in Fig. 3, for sufficiently large $a$, the wave speed is replaced by that corresponding to a compactly supported initial condition. This is because a trajectory may only end at a given $V_{\infty}$ when $V_{\infty}>\max\{V_{c},V_{s}\}$ for the given wave speed and value of $\gamma$. For sufficiently large $a$, the wave speed will be limited either by the existence of $V_{c}$, so that
$c=2(1-V_{\infty}),$ (14)
or by the existence of $V_{s}$, in which case $c$ will be the value such that $v_{0}=V_{s}$ for given $(\gamma,c)$.
@@
Using this procedure we produce curves of wave speed $c$ as a function of $\gamma$, for given values of the far-field resident cell population $V_{\infty}$. These results are plotted in Fig. 2 along with the estimates from the PDE simulations for compactly supported initial conditions. When $\gamma$ is sufficiently small such that $V_{s}$ does not exist, we instead use the condition that $V_{\infty}=V_{c}$, thus $c=2(1-V_{\infty})$. The comparison confirms that we have correctly identified the wave speed selection mechanism.
@@ 4 Large death rate asymptotic analysis
In this section we describe the asymptotic analysis of the relation between resident cell death rate $\gamma$ and wave speed $c$ in the limit that $\gamma$ becomes large. We restrict ourselves to the travelling wave branches that correspond to compactly supported initial conditions (that is, $V_{\infty}=V_{s}$ in the preceding section).
The important observation is that for large $\gamma$, (9b) implies that $V$ is asymptotically small except near the front (where $V\to V_{\infty}$ as $z\to\infty$). The effect of $\gamma$ is thus to modify the leading order, Fisher–KPP solution, a property that was observed in PDE simulations in El-Hachem et al. [11]. In addition, this phenomenon is very similar to that which occurs when the Fisher–KPP system is modified by a small cut-off factor in the reaction term (3), as considered in Brunet and Derrida [6] and Dumortier et al. [10]. Our asymptotic analysis will follow along similar lines to Brunet and Derrida [6]. We will also numerically determine the prefactors in the leading order asymptotic approximations, which are required to observe the effect of $V_{\infty}$ on the asymptotic wave speed.
@@ 4.1 Outer and intermediate regions
Let $\gamma\gg 1$ and let the wave speed $c=2-\delta$, where $\delta\to 0$ as $\gamma\to\infty$ is the correction to the wave speed. In the limit $\gamma\gg 1$, (9b) implies that $V=0$ to leading order, so that to leading order, $U$ and $W$ are given by travelling wave solutions to the Fisher–KPP system (1) with wavespeed $c=2$:
$U\sim U_{0}(z),\qquad W\sim W_{0}(z),\qquad\delta\to 0,\qquad z-z_{0}=O(1)$ (15)
@@
$U_{0}^{\prime}=W_{0},\qquad W_{0}^{\prime}=-2W_{0}-U_{0}(1-U_{0}),\qquad\lim_{z\to-\infty}(U_{0},W_{0})=(1,0).$ (16)
As $z\to\infty$, the behaviour of this leading order solution is determined by the repeated eigenvalue of the Fisher–KPP system near the origin:
$U_{0}\sim A_{0}(z-z_{0})\mathrm{e}^{-(z-z_{0})},\qquad z\to\infty.$ (17)
In (17) we have explicitly written the arbitrary constant of translation $z_{0}$. In this form the prefactor $A_{0}$ is not arbitrary, but is a constant that is a property of the unique solution to (16). The numerical value of $A_{0}$ may be found by computing the value of
@@
By solving (16) numerically and using the formula (18), we determine that
@@
As $z$ increases, the system (9) reaches an intermediate region in which the effect of subcritical wave speed ($\delta>0$) becomes important. Since this effect is to reduce the wave speed below the critical value $2$, the leading order Fisher–KPP approximation is pushed into the regime in which the origin is a stable spiral. That is,
$U\sim U_{M},\qquad(2-\delta)U_{M}^{\prime}+2U_{M}^{\prime\prime}+U_{M},$ (20)
with solution (matching to the outer solution (17))
$U_{M}=A_{0}\delta^{-1/2}\mathrm{e}^{-(z-z_{0})}\sin\left(\delta^{1/2}(z-z_{0})\right).$ (21)
This asymptotic behaviour holds when $\delta\to 0$ with $z-z_{0}=O(\delta^{-1/2})$. It is this intermediate behaviour that must be matched to the inner region in which $V$ is no longer small.

### v2

@@ 3 Calculation of travelling wave solutions
In this section we will compute travelling wave solutions to (8), and analyse their far field behaviour that determines which travelling wave solutions are selected from a given initial condition. Let z=xctz=x-ct be the travelling wave coordinate, and let $u(x,t)=U(z)$, $v(x,t)=V(z)$, and $\partial_{x}u(x,t)=W(z)$. The system of PDEs then becomes the system of three autonomous differential equations:
$\frac{\mathrm{d}U}{\mathrm{d}z}$ $=W,$ (11a)
$\frac{\mathrm{d}V}{\mathrm{d}z}$ $=\frac{\gamma}{c}UV,$ (11b)
$\frac{\mathrm{d}W}{\mathrm{d}z}$ $=\frac{1}{1-V}\left(\frac{\gamma}{c}UVW-cW-U(1-U-V)\right).$ (11c)
@@
The eigenvalues and eigenvectors of the fixed point $(1,0,0)$ are
@@
$\lambda_{1}^{\prime}=0,\quad\lambda_{2}^{\prime}=-\mathcal{C}+\sqrt{\mathcal{C}^{2}-1},\quad\lambda_{3}^{\prime}=-\mathcal{C}-\sqrt{\mathcal{C}^{2}-1},\quad\mathcal{C}=\frac{c}{2(1-V_{\infty})},$ (13)
@@
The zero eigenvalue $\lambda_{1}^{\prime}$ (with eigenvector pointing along the $V$-axis) is due to the non-isolated nature of each fixed point. The other two eigenvalues are negative, indicating that trajectories are attracted to the $V$-axis in its neighbourhood. Because the unstable manifold is two-dimensional, there is one degree of freedom in trajectories leaving $(1,0,0)$. There is thus a one-parameter family of trajectories that connect $(1,0,0)$ to the $V$-axis, each ending at a different value of $V_{\infty}$ between 0 and $1$.
We compute solutions of (11) numerically by solving it as a boundary value problem on a large domain $z\in(0,z_{\infty})$, using the numerical continuation package Auto-07p [8]. The use of numerical continuation will allow us to compute solutions for extreme values of parameters (in particular, very large $\gamma$). Appropriate boundary conditions are
$U(0)=1+\varepsilon,\quad W(0)=-\varepsilon\lambda_{2},\quad V(z_{\infty})=V_{\infty},$
where $\varepsilon$ is a small parameter and $z_{\infty}\gg 1$; in our calculations we take $\varepsilon=10^{-5}$ and $z_{\infty}=50$. These boundary conditions enforce that near $(1,0,0)$, trajectories are on the eigenvector with the smaller positive eigenvalue. For a given $\gamma$ and $c$ we are able to construct travelling wave solutions for any value of $V_{\infty}$ between 0 and 1, by starting at $V_{\infty}=0$ (at which point $V$ vanishes identically, and $U,W$ are solutions to the Fisher–KPP system), and then increasing $V_{\infty}$ by numerical continuation.
We plot examples of such trajectories, projected onto the $(U,V)$ plane, in Fig. 4. In this figure we choose two sets of parameters $(\gamma,c)=(1,1)$ and $(10,1.24)$ in order to compare to the travelling wave solutions observed in the PDE simulations shown in Fig. 1. For given values of $\gamma$ and $c$, there is a one-parameter family of travelling wave solutions, that we may consider as characterised by the value $V_{\infty}$. Since the wave speed $c$ is a parameter of the system, we thus expect that for a given $V_{\infty}$ there is a one-parameter family of travelling wave solutions, parametrised by $c$. In Fig. 4 we observe that the travelling waves that emerge from the PDE simulations do indeed correspond to certain trajectories in phase space. However, we are yet to determine which trajectory (in particular, which wave speed $c$) will correspond to the long-time behaviour of a solution to the PDE system (8) for a given initial condition. In order to determine this wave-speed selection, we need to consider the behaviour of trajectories near the front of the travelling wave, that is, near the $V$-axis.
@@
For initial conditions of the form $u(x,0)\sim\mathrm{e}^{-ax}$, $x\to\infty$, as shown in the previous section, one may expect the appropriate travelling wave to be that which has the same exponential decay. Since this decay is given by the least negative eigenvalue $\lambda_{2}^{\prime}$ in (13), by equating $a=-\lambda_{2}^{\prime}$ we find the relationship
$c=(1-V_{\infty})\left(a+\frac{1}{a}\right).$ (14)
For initial conditions that have sufficiently slow decay this relationship does indeed hold. For example, the solutions shown in Fig. 1(c) and Fig. 1(d), corresponding to values $(v_{0},\gamma,c)=(0.75,1,1)$, and $(v_{0},\gamma,c)=(0.75,10,1.24)$, satisfy this relationship. In the phase plane in Fig. 4 it is also observed that these PDE simulations correspond to trajectories in phase space that terminate at $V_{\infty}=0.75$ for the appropriate values of $\gamma$ and $c$. This relationship can also be seen in the results of simulations depicted in Fig. 3; for values of $a$ such that $c>2$, that is,
$a<\frac{1}{1-V_{\infty}}-\sqrt{\frac{1}{(1-V_{\infty})^{2}}-1},$ (15)
the wave speed is always determined by the criterion (14). For larger $a$, this relation holds for sufficiently small $\gamma$, until a greater constraint on the wave speed is encountered.
As occurs in the analysis of the Fisher–KPP system, a further criterion results from avoiding parameters that result in travelling waves in which $U$ does not remain positive over all $z$. This will be important for sufficiently large decay rate $a$, and in particular compactly supported initial conditions. Firstly, we note that for a given wavespeed $c$, there is a critical value $V=V_{c}$, at which the eigenvalues $\lambda^{\prime}_{2,3}$ switch from complex to real:
@@
This value corresponds to $\mathcal{C}=1$ in (13). Notably, $V_{c}$ is independent of the decay rate $\gamma$. Trajectories for which $V_{\infty}<V_{c}$ spiral into the relevant point on the $V$-axis, and since $U$ will become negative on such a spiral, these trajectories cannot correspond to solutions to (8) that have evolved from a non-negative initial condition. This property is determined locally, and is equivalent to the condition used to determine the minimum wave speed in the Fisher–KPP equation. It results in the prediction for the wave speed:
$c=2(1-V_{\infty}),\qquad V_{\infty}=V_{c}.$ (17)
This relation can also be observed in the PDE simulations; it is the relevant constraint for the solution $v_{0}=0.5,\gamma=1,c=1$ as depicted in Fig. 1(a), and as depicted by the dashed lines in Fig. 2, for sufficiently small $\gamma$.
However, (17) does not explain the dependence of wave speed on $\gamma$ when $\gamma$ is sufficiently large, so the wave speed in this regime must be limited by a different criterion. The reason for this additional constraint is that for some parameter regimes, $U$ becomes negative as $z\to\infty$, even when $V_{\infty}>V_{c}$, that is, when the eigenvalues of the steady state are real. In Fig. 4(b) (in particular the inset of that figure), in which $c=1.24$ and $\gamma=10$, we observe that trajectories for which $V_{\infty}$ is slightly larger than $V_{c}$ still approach the $V$-axis from the negative $(U<0)$ side, despite the fact that eigenvalues are real at these points. In this parameter regime there is a point $V_{s}$ (indicated as a red dot in Fig. 4(b)) that separates trajectories that have $U$ become negative, from those that remain positive. Also seen in this figure is that this is the appropriate trajectory corresponding to the PDE simulation depicted in Fig. 1(b), where $(v_{0},\gamma,c)=(0.5,10,1.24)$, which does not satisfy the condition (17). The existence of such a threshold value was also noted in the similar model of Colson et al. [7] (7) described in the introduction. Unlike the point $V_{c}$, the point $V_{s}$ is not determined as a local property of the phase space (it is a property of the manner in which the unstable manifold of $(1,0,0)$ intersects the $V$-axis), so is unlikely to have a closed-form expression.
@@
Using this procedure we produce curves of wave speed $c$ as a function of $\gamma$, for given values of the far-field resident cell population $V_{\infty}$. These results are plotted in Fig. 2 along with the estimates from the PDE simulations for compactly supported initial conditions. When $\gamma$ is sufficiently small such that $V_{s}$ does not exist, we instead use the condition (17) assuming $V_{\infty}=V_{c}$. The comparison confirms that we have correctly identified the wave speed selection mechanism.
@@ 4 Large death rate asymptotic analysis
In this section we describe the asymptotic analysis of the relation between resident cell death rate $\gamma$ and wave speed $c$ in the limit that $\gamma$ becomes large. We restrict ourselves to the travelling wave branches that correspond to compactly supported initial conditions (that is, $V_{\infty}=V_{s}$ in the preceding section). We introduce the wave speed correction $\delta$, such that the wave speed is $c=2-\delta$. From the above numerical results, $\delta\to 0$ as $\gamma\to\infty$, and the main task of the asymptotic analysis is to determine how $\delta$ depends on $\gamma$ in this limit.
The asymptotic structure that arises in the large $\gamma$ limit is depicted in Fig. 6. The important observation is that for large $\gamma$, (11b) implies that $V$ is exponentially small except near the front (where $V\to V_{\infty}$ as $z\to\infty$). There are thus three main regions: an outer region, where to leading order $U$ is given by the critical wavespeed ($c=2$) solution of the Fisher–KPP equation; an intermediate region, where $V$ is still negligible but the effect of slightly subcritical wavespeed is felt (that is, $\delta>0$); and finally an inner region, where $V$ must be order unity to match with the far field condtion ($V\to V_{\infty}$), and $U=O(\gamma^{-1})$ is small. These regions are named in reference to solutions in phase space, where the outer region is the part of the phase space close to the $(U,W)$ plane, while the inner region corresponds to the neighbourhood of the $V$-axis, on which $(U,W)=(0,0)$.
The asymptotic structure of our system is very similar to that which occurs when the Fisher–KPP system is modified by a small cut-off factor in the reaction term (3), as considered in Brunet and Derrida [3] and Dumortier et al. [9]. In our case, $\gamma^{-1}$ plays the role of the small parameter $\epsilon$ of that model. The relationship between $\gamma$ and the wavespeed correction $\delta$ is found by matching the solutions in the intermediate and inner regions. To leading order this is universal (indeed, the same as (4)) and does not depend on the value of $V_{\infty}$. However, we will also numerically determine the prefactors in the leading order asymptotic approximations, which are required to observe the effect of $V_{\infty}$ on the asymptotic wave speed.
Figure 6: The asymptotic structure of travelling wave solutions in the large death rate ($\gamma\to\infty$) limit. The wave speed $c=2-\delta$ where $\delta\to 0$ as $\gamma\to\infty$. In the outer region, $V$ is exponentially small, and $(U,W)$ is given by the Fisher-KPP travelling wave with $c=2$. In the intermediate region, the effect of sub-critical wave speed $(0<\delta\ll 1)$ is felt, while $V$ remains negligible. In the inner region, the effect of the far field condition enforces $V=O(1)$. The dependence of $\delta$ on $\gamma$ is found by matching between the intermediate and inner regions.
@@ 4.1 Outer and intermediate regions
We first consider the outer region depicted in Fig. 6. In the limit $\gamma\gg 1$, (11b) implies that $V$ is exponentially small, so that to leading order, $U$ and $W$ are given by travelling wave solutions to the Fisher–KPP system (1) with wavespeed $c=2$:
$U\sim U_{0}(z),\qquad W\sim W_{0}(z),\qquad\delta\to 0,$ (19)
@@
$\frac{\mathrm{d}U_{0}}{\mathrm{d}z}=W_{0},\qquad\frac{\mathrm{d}W_{0}}{\mathrm{d}z}=-2W_{0}-U_{0}(1-U_{0}),\qquad\lim_{z\to-\infty}(U_{0},W_{0})=(1,0).$ (20)
As it will be required to compute prefactors, we have numerically calculated this solution, using the standard method of solving the initial value problem starting on the unstable eigenvector near $U_{0}=1,W_{0}=0$. A plot of $U_{0}(z)$ is shown in Fig. 7(a).
Figure 7: (a) Numerically computed solution $U_{0}(z)$ of the leading-order outer problem (20), which is the Fisher-KPP system. (b) Plot of the expression $\ell(z)$ (22) determined from $U_{0}$ and $W_{0}=U_{0}^{\prime}$, which is used to find the prefactor $A_{0}$ in the $z\to\infty$ limit (21) of the leading order problem. (c) Fit of the far field behaviour (21) with the numerical solution (plotted on a logarithmic scale), showing that when $z$ is equal to the reference value $z_{0}$, $U_{0}\approx 0.149$.
As $z\to\infty$, the behaviour of this leading order solution is determined by the repeated eigenvalue with value 1-1 of the Fisher–KPP system (20) near the origin. This repeated eigenvalue implies that to leading order,
$U_{0}\sim Az\mathrm{e}^{-z},\qquad z\to\infty$
where the prefactor $A$ is arbitrary, as the system (20) is translationally invariant in $z$. However, in order to carry out a matched asymptotic expansion to the order required to observe the effect of different values of $V_{\infty}$, we require a more precise form of the far field behaviour of $U_{0}$. In particular, we write
$U_{0}=A_{0}(z-z_{0})\mathrm{e}^{-(z-z_{0})}+o\left(\mathrm{e}^{-z}\right),\qquad z\to\infty,$ (21)
where we have explicitly included the translation constant $z_{0}$. In this form, the prefactor $A_{0}$ is not arbitrary, but is a constant that is fixed by stipulating that the order of error be less than exponential. Since $A_{0}$ depends on the trajectory of the solution as a whole, it can only be estimated from a numerical solution of (20).
One way to estimate the value of $A_{0}$ from such a numerical solution of (20) is to differentiate (21) to find $W_{0}$, and eliminate $A_{0}$ to find
$U_{0}+W_{0}\sim A_{0}\mathrm{e}^{-(z-z_{0})},\qquad\frac{U_{0}}{U_{0}+W_{0}}\sim z-z_{0},\qquad z\to\infty.$
By taking the logarithm of the first of these, and using the second to eliminate $z-z_{0}$, we find
@@
In Fig. 7(b) we plot $\ell(z)$ over time, and observe that it tends to a constant $\approx-1.953$ as $z\to\infty$. Using this we determine that the prefactor is
@@
In writing (21), the constant of the translation $z_{0}$ may be thought of as a reference point in the outer region, corresponding to the point where $U_{0}$ takes a specific (order one) value. Indeed, this value may also be estimated from the numerical solution by fitting (21), with our computed value of $A_{0}$, to the numerical solution. We show this fit in Fig. 7(c). While the numerical value of $z_{0}$ is arbitrary, as the problem is translationally invariant, we find that it corresponds to the point where $U_{0}(z_{0})\approx 0.149$. In terms of the asymptotic structure of the problem, what is important is that $z_{0}$ is a point in the outer region where $U=O(1)$ is a fixed value.
We now consider the intermediate region in which the effect of subcritical wave speed ($\delta>0$) becomes important while $V$ is still exponentially small. As the wave speed is below the critical value $2$, the leading order Fisher–KPP approximation is pushed into the regime in which the origin is a stable spiral. That is,
$U\sim U_{M},\qquad(2-\delta)\frac{\mathrm{d}U_{M}}{\mathrm{d}z}+\frac{\mathrm{d}^{2}U_{M}}{\mathrm{d}z^{2}}+U_{M}=0,$ (24)
with asymptotic solution (matching to the outer solution (21))
$U_{M}\sim A_{0}\delta^{-1/2}\mathrm{e}^{-(z-z_{0})}\sin\left(\delta^{1/2}(z-z_{0})\right).$ (25)
Here the prefactor $A_{0}\delta^{-1/2}$ results from matching with the outer solution (19), so that expanding $U_{M}$ in small $\delta$ results in the leading order behaviour (21) as $z\to\infty$. This asymptotic behaviour holds when $\delta\to 0$ with $z-z_{0}=O(\delta^{-1/2})$, thus the magnitude of $U_{M}$ is exponentially small despite the algebraically large prefactor.
The intermediate solution is valid until the sinusoidal term in (25) approaches zero. In this limit, the intermediate solution must be matched to the inner solution where $V$ is no longer negligible. Noting that $z=z_{0}$ is not in the intermediate region, the first (that is, the smallest) value of $z$ at which this occurs is when $z=z_{0}+\delta^{-1/2}\pi$. The matching between the intermediate and inner problems is where the dependence of $\delta$ on $\gamma$ will be determined, as we describe next.